OpenStudy (anonymous):

If a/4 = 9/a, then...? 1. a²=36 2. 4a = 9a 3. a + 4 = 9 + a 4. a - 4 = 9 - a

OpenStudy (studygurl14):

Cross multiply. \(\large\frac{a}{4}=\frac{9}{a}\rightarrow aa = 4(9)\rightarrow ??\)

OpenStudy (anonymous):

Thanks.